OpenStudy (anonymous):

Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v. a. <-20, -42> b. <-20, 0> c. <0, 18> d. <-6, 0>

OpenStudy (perl):

-2u +5v = -2 < 5,6> + 5<-2,-6> = <-2*5, -2*6 > + <5*-2, 5*-6> = <-10,-12> + <-10,-30> = <-10 + (-10) , -12 + (-30) > = <-20, -42> please check my work

OpenStudy (anonymous):

okay you got it i see what i was doing wrong now:) thxx